OpenStudy (anonymous):

what is the equation of a polynomial function that has zeros at -4,-2, and 1. and that passes through the point (-3,12)

OpenStudy (anonymous):

\[x^3+5 x^2+2 x-8\]

OpenStudy (amistre64):

well, if we use whats given and construct one with what we should know; it will tend to just coaless into the proper form

OpenStudy (amistre64):

it has 3 zeros; meaning 0*0*0 = 0 (x+4)(x+2)(x-1)=0 now to antifoil lol (x+4) (x+2) ----- x^2 +4x +2x +8 ------------ x^2 +6x +8 x-1 ------------ x^3 +6x^2 +8x -x^2 -6x -8 ---------------- x^2 +5x^2 +2x -8 x^2 +5x^2 +2x -8; is the base function, which we need to modify to hit the given point, so lets start by seeing how far off we are

OpenStudy (amistre64):

when (x=-3,y=12) x^3 +5x^2 +2x -8 = y (-3)^3 +5(-3)^2 +2(-3) -8 = 12 -27 +45 -6 -8 = 12 -27 +45 -14 = 12 -41 +45 = 12 4 = 12 ; the only modification we can make is a scaled modification, in other words multiply or divde 4s = 12 when s = 3 y = 3(x^3 +5x^2 +2x -8) y = 3x^3 +15x^2 +6x -24

OpenStudy (jamesj):

Or in other words DHASHNI was wrong.

OpenStudy (amistre64):

Dhashni provide the base model :)

OpenStudy (amistre64):

we had to add the customized headers and pleather front bench seating

OpenStudy (anonymous):

thanks man, helped alot

OpenStudy (jamesj):

A slightly more compact way of dealing with 'passes through (-3,12) is as follows. The polynominal must have the form for some number A p(x) = A(x+4)(x+2)(x-1), as -4, -2 and 1 are roots. So if (-3,12) lies on the graph, p(-3) = 12 I.e., A.1.-1.-4 = 12, hence 4A = 12, A = 3 So the polynominal is p(x) = 3(x+4)(x+1)(x-1)

OpenStudy (anonymous):

ANTIFOIL!

OpenStudy (amistre64):

.... whenever i do math my knuckles bleed :)